OpenStudy (anonymous):

Ok I'm confused. The question is... Find the area of a circle if the radius is 5 meters. Leave the answer in terms of pi. I keep getting 78.5 but that's not an answer choice. Im doing the correct steps so whats wrong???

OpenStudy (anonymous):

A= pi*r^2

OpenStudy (anonymous):

19.625, try the formula and see if you get the same.

OpenStudy (anonymous):

@Miloo, are you trying?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

2.5*25.=6.25 6.25*3.14=19.625

OpenStudy (anonymous):

Ok I was multiplying wrong with the formula I see how you got the answer to the formula.

OpenStudy (anonymous):

is it an answer choice

OpenStudy (anonymous):

No the answer choices are ) A =100π square meters B) A =25π square meters Eliminate C) A = 5π square meters D) A = 10π square meters

OpenStudy (anonymous):

Where did you get 2.5 I thought pi was 3.14?

OpenStudy (anonymous):

i was squaring thr radius first

OpenStudy (anonymous):

This is the only picture is gives me showing the radius and diameter.

OpenStudy (triciaal):

why did you eliminate the correct answer?

OpenStudy (triciaal):

Area A = pi r^2 r = 5 leave pi as pi substitute A = pi (5^2) A = 25 pi unit of measure m x m = meter square

OpenStudy (anonymous):

I didnt eliminate it. When I copied it the eliminate button showed up. Sorry but I see how you got the answer. Thanks for everyones help.

OpenStudy (triciaal):

welcome